# Symmetry-breaking induced surface magnetization in non-magnetic RuO2

### v1

@@ Abstract
Altermagnetism is a newly identified phase of magnetism distinct from ferromagnetism and antiferromagnetism. RuO2 has been considered a prototypical metallic altermagnet with a critical temperature higher than room temperature. Previous interpretations of unusual magnetic properties of RuO2 relied on the theoretical prediction that local moments on two Ru sublattices, which are connected by four-fold rotational symmetry, are quite significant ($\sim 1\,\mu_{B}$), leading to long-range antiferromagnetic order. However, accumulated experimental data suggest that local moments on Ru in RuO2 are vanishingly small, indicating the bulk material is likely non-magnetic. This observation is consistent with the delocalized nature of the $4d$ electron of Ru and the strong screening effect in the metallic state. In this work, we show that despite the non-magnetic bulk, RuO2 surface (110) exhibits spontaneous magnetization, which we attribute to the breaking of local symmetry, resulting in electronic redistribution and magnetic moment enhancement. The emergence of surface magnetism gives rise to interesting spectroscopic phenomena, including spin-polarized surface states, spin-polarized scanning probe microscopy images, and potentially spin-dependent transport effects. These highlight the important role of surface magnetic structures in the otherwise non-magnetic bulk RuO2.
@@ I Introduction
Altermagnetism (AM) is an emerging phase of magnetism that differs from the well-known ferromagnetism (FM) and antiferromagnetism (AFM)  [1]. In altermagnetic materials, spin-splitting phenomena of the electronic band structure occur alternately in reciprocal space directions that are related to each other by rotation or reflection but not translation or inversion, which is determined by the presence of non-magnetic ligands in real space [1]. Among the earliest and most studied candidates for altermagnetic materials is RuO2, which has become a prototype in this field [1, 2]. Despite various theoretically predicted [2] and experimentally observed phenomena such as the anomalous Hall effect [3], time-reversal symmetry breaking [4], ascribed to the assumption that RuO2 is an antiferromagnet with a high Néel temperature and a significant local moment (typically assumed $\sim 1\,\mu_{B}$ in theoretical calculations), direct experimental evidence for such a high moment in this $4d$ metallic oxide is lacking. Therefore, the magnetic configuration of RuO2, especially the presence of local moments and long-range magnetic ordering, is still under active debate. The first neutron scattering experiment on RuO2 reported a local moment of only $0.05\,\mu_{B}$ [5]. Subsequently, an experiment using the resonant X-ray scattering technique supported the ground state of AFM in RuO2 [6]. However, recent in-depth investigations strongly suggest that the ground state of RuO2 is non-magnetic/paramagnetic [7, 8, 9, 10, 11]. For instance, highly local-moment-sensitive muon spin rotation ($\mu$SR) and neutron diffraction experiments have shown that the ground state of the bulk RuO2 is likely non-magnetic, with local moments on the order of $10^{-4}\,\mu_{B}$ [7, 9]. Furthermore, by using broadband infrared spectroscopy combined with first-principles calculations to probe the optical conductivity of RuO2—its bulk electronic property—researchers concluded that bulk RuO2 is well described by a non-magnetic model [11]. Very recently, Liu et al. have directly probed the electronic structure of RuO2 in bulk and thin film samples using angle resolved photoemission spectroscopy (S-ARPES) and found no evidence of spin splitting in the electronic structure of the material [8]. These findings align with the delocalization and weak correlation nature of the Ru $4d$ orbitals, consistent with the metallic property of the compound.
@@
Most of the experiments mentioned above have been concentrated on bulk or thick samples, with less focus on surface properties. However, the surface properties of altermagnets might play a significant role in various phenomena at surfaces and interfaces, or even be the key to explaining some observations [12, 13]. Of many low-index surfaces, RuO2’s surface and thin films in the crystallographic direction $\langle 110\rangle$ are of particular interest since that is the crystal surface with the lowest surface energy and is naturally cleaved [14, 15]. For example, the RuO2 (110) surface has been a primary focus in catalysis research thanks to its high electrocatalytic water splitting activity, attributed to the presence of Ru atoms with dangling bonds [14, 16]. However, most studies have ignored the magnetic properties of the surface, except for a couple of theoretical predictions, which highlighted the role of surface magnetism in RuO2 (110) in catalytic reactions during electrolysis [17, 18]. Interestingly, strained RuO2 (110) thin films grown on TiO2 substrates have recently been shown to exhibit superconducting behavior at low temperatures and a metal-to-insulator transition in the ultrathin limit regime [19, 20, 21]. The role of the soft-phonon mode and the enhancement of electron density at the Fermi level ($E_{F}$) have been invoked to explain this superconductivity. However, the possible role of the magnetic structure in supporting the superconductivity of these thin films has been overlooked.
@@
As noted previously, Torun et al. [17] have theoretically predicted that surface magnetization can spontaneously develop on the RuO2 (110) surface. However, their study was not without limitations. First, the slab they used in their calculations, consisting of only five layers of RuO2, was probably insufficient to accurately represent both surface and bulk-like regions, potentially affecting the predicted behavior of surface magnetization. In metallic systems such as RuO2, the electronic wavefunctions of the top and bottom surfaces of the slab used in the calculation can penetrate deep into the bulk and mutually interact if the thickness of the slab is not sufficient, leading to an alteration of electronic properties of the true surface. Second, the origin of the surface magnetism was not comprehensively explored. More importantly, since their work focused on surface catalysis, it did not address how surface magnetization influences electronic and magnetic properties that are critical for spectroscopic and transport phenomena, which is highly relevant for spintronics applications [4, 22].
@@
In this paper, by using first-principles calculations employing Density Functional Theory (DFT), we show that surface magnetism can spontaneously be developed in RuO2 (110). The spontaneous surface magnetization can be understood from the breaking of symmetry on the surface of RuO2 (110), leading to a significant reconstruction of the electronic structure of RuO6 octahedra at the surface layer, i.e., in the band filling of the Ru $4d$ orbital, resulting in the emergence of sizeable local moments. Remarkably, the presence of spontaneous surface magnetization leads to the emergence of spin-polarized surface states and spin-dependent transport effects. These results can be important for understanding the experimental data on RuO2 obtained using spectroscopic and transport measurements such as spin- and angle-resolved photoemission spectroscopy (S-ARPES) [23], spin-polarized scanning tunneling microscopy/spectroscopy (SP-STM/SP-STS), anomalous Hall measurements, and interfacial spin-dependent phenomena.
@@ II Computational Method
Bulk RuO2 was simulated using a tetragonal unit cell with a rutile structure ($P4_{2}/mnm$, SG 136). Due to the metallic property of bulk RuO2, a dense $k$-point mesh of 12×12×18 was used to sample the bulk BZ. The lattice constant was adopted from a recent experimental determination [32]. RuO2 films along [110] were built by reorienting the lattice vectors of the primitive cell, i.e., primitive lattice vectors along [001], [$01\overline{1}$ 0], and [110] become [100], [010], and [001] of the RuO2 (110) surface unit cell, respectively, thus giving us a nonpolar stoichiometric slab. The symmetry of the bulk unit cell transitions to $Pmmm$ (SG 47) for stoichiometric slabs with an odd number of Ru layers. A $k$-point mesh of 12×6×1 was used for structural optimization of the surface unit cell, and a denser mesh of 18×9×1 was used for density of state (DOS) calculations. To eliminate spurious image interactions between slabs, a vacuum space of at least 20 Å was added to the slab normal direction. Internal coordinates of the slab were fully relaxed until the Hellman-Feynman force on each atom was smaller than 0.005 eV/Å while keeping the lattice vectors fixed to the experimental bulk values. Due to the recent experimental observation of a non-magnetic ground state in bulk RuO2 [7, 9], we employed the PBE+$U$ approach with $U=0$ [33, 34, 35] to describe its electronic structure. Additionally, since spin-orbit coupling does not significantly affect the electronic and magnetic structures of RuO2, it was neglected in our calculations.
@@
To gain insights into the chemical bonding between atomic pairs contributing to the emergence of spin-polarized states at the surface of RuO2, we performed a crystal orbital Hamilton population (COHP) analysis [36, 37]. The COHP method decomposes the one-particle band energies into interactions between atomic orbitals of adjacent atoms. It effectively weighs the electronic density of states (DOS) by the corresponding Hamiltonian matrix elements, thus recovering the phase information of calculated wavefunctions that is otherwise missing in the band structure or DOS descriptions of material electronic structures. This enables the identification of bonding, non-bonding, and antibonding characteristics among the pairwise atoms. The stability of these interactions is quantified by the COHP values. A positive COHP value indicates bonding interactions, while a negative value represents antibonding character. Traditionally, COHP analysis was done directly from atom-centered basis set calculations. However, since our calculations were done using the plane-wave basis set, a projection scheme from this basis set to a local orbital basis set was employed using the LOBSTER code [36, 37, 38]. In this projection scheme, the equivalence of the traditional COHP quantity is --pCOHP (negative of projected COHP). In our study, we applied COHP analysis to examine the nearest-neighboring atomic pairs around Ru atoms both on the surface and in bulk-like regions. For the projection, we employed the local basis function as defined in pbeVaspFit2015.
@@
Since spin-polarized scanning tunneling microscopy (SP-STM) technique can be used to recognize spin polarization effects on the surface [39], we simulated spin-resolved STM images employing the Tersoff-Hamann approximation [40]. In this approach, the tunneling current at the simulated probe tip (i.e., at a particular distance from the surface) in an STM experiment is proportional to the local density of states (LDOS) of the integrated electronic states ranging from the Fermi level to a predefined energy level given by a bias voltage $V$, roughly corresponding to an experimental voltage. The LDOS is given by
$n(r,E)=\sum_{\mu}\left|\psi_{\mu}(r)\right|^{2}\delta\left(\varepsilon_{\mu}-E\right),$ (1)
and the tunneling current is expressed as
@@
where $n(r,E)$ represents the LDOS at a given position $r$ and energy $E$. The LDOS can be evaluated from partial charge densities calculated from a pre-converged wavefunction. The partial charge density file from VASP was read by the HIVE-STM program [41], and STM images were generated using the constant-height method (at a height of $\sim 3$ above the highest atoms of the slab—the oxygen bridging atoms). For a given bias $V$, the simulated STM images reflect the contrast in the partial charge densities within the energy range $0<E-E_{F}<V$ (for $V>0$, positive bias) or $V<E-E_{F}<0$ (for $V<0$, negative bias).
@@ III.1 Electronic and magnetic properties of bulk RuO2
Figure 1: Rutile-type crystal structure, electronic structure and chemical bonding analysis of bulk RuO2. (a) Rutile structure of RuO2 with two Ru sublattices in red and blue, oxygen in silver color, (b) Bulk Brillouin zone (BZ) shown in black and rotated so that [110] crystallographic direction normal to the (110) surface whose BZ shown in red, (c) spin-polarized electronic band structure of the bulk exhibiting no spin-polarization (Fermi level set to 0), (d) atomic projected density of state for Ru and O in bulk, and (e) the corresponding --pCOHP curve on the same energy scale.
The spin-resolved electronic band structure of the bulk is shown in Fig. 1(c), which does not exhibit spin polarization. In the assumed altermagnetic state, the electronic structure of interest lying along the $\Gamma M$ high-symmetry path of the bulk Brillouin zone (BZ) shows significant spin splitting ($>1$ eV) as described by theoretical calculations of the band structure with relatively large on-site Hubbard $U$ values ($U\sim 1.5\div 2$ eV) [2, 5]. However, this spin splitting does not appear in the non-magnetic state. Along this momentum direction, three bands are present near the Fermi level, with an additional flat band lying around $-1.5$ eV below the Fermi level. When considering the electronic structure of the (110) slab, the $\Gamma M$ path of the bulk BZ is of particular importance since bulk band dispersions along this path and parallel to it are projected onto the high-symmetry path $\overline{\Gamma}\overline{Y}$ of the surface BZ, as illustrated in Fig. 1(b).
The density of state plot and chemical bonding analysis, shown in Fig. 1(d) and  1(e), respectively, highlights important characteristics of Ru–O bonds in RuO2. The DOS plot in Fig. 1(d) clearly shows strong hybridization between the Ru and O orbitals over a wide range of energy. Notably, between 8 and 6 eV below the Fermi level, pronounced peaks in the atomic-projected density of states (PDOS) and --pCOHP (positive—bonding character) with similar contributions from both Ru and O indicate a strong covalent bonding, best described by the $\sigma$-bonding orbitals formed by Ru $4d$ ($e_{g}$) and O $2p$ components [42]. In addition, approximately from 6 eV to 3 eV below the Fermi level, small positive values of --pCOHP, with a greater contribution of oxygen, suggest a weaker covalent bonding character of the Ru–O bonds due to the $\pi$-bonding between Ru $4d$ ($t_{2g}$) and O $2p$. Around $-2$ eV, the contribution of Ru disappears, i.e. Ru PDOS is zero, leaving only O $2p$ states. This indicates that the electrons occupying these states belong to non-bonding orbitals and do not contribute to the Ru–O bonding strength, leading the --pCOHP values to approach zero.
@@
Despite the non-magnetic character of bulk RuO2, intriguing phenomena can arise at the (110) surface, leading to the emergence of surface magnetism. This disparity between bulk and surface properties can be attributed to the altered electronic structure and bonding characteristics at the surface, which disrupt the symmetry and electronic interactions present in the bulk. The unique environment at the surface can facilitate localized magnetic moments that do not exist in the bulk material. In the next section, we will explore how these localized states, influenced by surface geometry and reduced coordination, contribute to the overall magnetic behavior of the (110) surface.
III.2 Surface magnetization in RuO2 (110) surface
III.2.1 The emergence of surface magnetization in RuO2 (110) surface

### v2

@@ Abstract
Altermagnetism is a newly identified phase of magnetism distinct from ferromagnetism and antiferromagnetism. RuO2 has been considered a prototypical metallic altermagnet with a critical temperature higher than room temperature. Previous interpretations of unusual magnetic properties of RuO2 relied on the theoretical prediction that local moments on two Ru sublattices, which are connected by four-fold rotational symmetry, are quite significant ($\sim 1\,\mu_{B}$), leading to long-range antiferromagnetic order. However, accumulated experimental data suggest that local moments on Ru in RuO2 are vanishingly small, indicating the bulk material is likely non-magnetic. This observation is consistent with the delocalized nature of the $4d$ electron of Ru and the strong screening effect in the metallic state. In this work, we show that despite the non-magnetic ground state in the bulk, RuO2(110) surface exhibits spontaneous magnetization, which we attribute to the breaking of local symmetry, resulting in electronic redistribution and magnetic moment enhancement. The emergence of surface magnetism gives rise to interesting spectroscopic phenomena, including spin-polarized surface states, spin-polarized scanning probe microscopy images, and potentially spin-dependent transport effects. These highlight the important role of surface magnetic structures in the otherwise non-magnetic bulk RuO2.
@@ I Introduction
Altermagnetism (AM) is an emerging phase of magnetism that differs from the well-known ferromagnetism (FM) and antiferromagnetism (AFM)  [1]. In altermagnetic materials, spin-splitting in the electronic band structure occurs alternately in reciprocal space directions that are related to each other by rotation or reflection but not translation or inversion, which is determined by the presence of non-magnetic ligands in real space [1]. Among the earliest and most studied candidates for altermagnetic materials is RuO2, which has been considered a prototype altermagnet [1, 2]. Despite various theoretically predicted [2] and experimentally observed phenomena linked to time-reversal symmetry breaking in AM such as the anomalous Hall effect [3], spin-splitter effect [4], magnetic circular dichroism [5], ascribed to the assumption that RuO2 is an AM with a high Néel temperature and a significant local moment (typically assumed $\sim 1\,\mu_{B}$ in theoretical calculations), direct experimental evidence for such a high moment in this $4d$ metallic oxide is lacking. Therefore, the magnetic configuration of RuO2 and especially the presence of local moments and long-range magnetic ordering are still subjects of debate [6]. The first neutron scattering experiment on RuO2 reported a local moment of only $0.05\,\mu_{B}$ [7]. Subsequently, an experiment using the resonant X-ray scattering technique supported the AFM ground state in RuO2 [8]. However, recent in-depth investigations strongly suggest that the ground state of RuO2 is non-magnetic/paramagnetic [9, 10, 11, 12, 13, 14]. For instance, local-moment-sensitive muon spin rotation ($\mu$SR) and neutron diffraction experiments have shown that the ground state of the bulk RuO2 is likely non-magnetic, with negligibly small local moments of $10^{-4}\,\mu_{B}$ [9, 11]. Furthermore, by using broadband infrared spectroscopy combined with first-principles calculations to probe the optical conductivity of RuO2—a bulk electronic property—researchers concluded that bulk RuO2 is well described by a non-magnetic model [13]. Very recently, Liu et al. have directly probed the electronic structure of RuO2 in bulk and thin film samples using angle-resolved photoemission spectroscopy (S-ARPES) and found no evidence of spin splitting in the electronic structure of the material [10]. These findings align with the delocalization and weak correlation nature of the Ru $4d$ orbitals, consistent with the metallic property of the compound.
@@
Most of the experiments mentioned above have been concentrated on bulk or thick films, with less focus on surface properties. However, the surface properties of altermagnets might play a significant role in various phenomena at surfaces and interfaces or even be key to explaining some observations [15, 16]. Of many low-index surfaces, RuO2 surface and thin films in the crystallographic direction $\langle 110\rangle$ are of particular interest since that is the crystal surface with the lowest surface energy and is of easy cleavage [17, 18]. For example, the RuO2(110) surface has been a primary focus in catalysis research thanks to its high electrocatalytic water splitting activity, attributed to the presence of Ru dangling bonds [17, 19]. However, most studies have ignored the magnetic properties of the surface, except for a couple of theoretical predictions, which highlighted the role of surface magnetism in RuO2(110) in catalytic reactions during electrolysis [20, 21]. Given the subtlety of magnetism in RuO2 seen in the literature, it is possible that surface effect could be a relevant source of magnetic phenomena observed in various samples. For instance, strained RuO2(110) thin films grown on TiO2 substrates have recently been shown to exhibit superconducting behavior at low temperatures and a metal-to-insulator transition in the ultrathin limit regime [22, 23, 24]. The role of the soft-phonon mode and the enhancement of electron density at the Fermi level ($E_{\text{F}}$) have been invoked to explain this superconductivity. However, the possible role of the magnetic structure in supporting these phenomena has been overlooked.
@@
As noted previously, Torun et al. [20] have predicted that surface magnetization can spontaneously develop on the RuO2(110) surface. However, their study was not without limitations. First, the slab used in their calculations, consisting of only five layers of RuO2, was probably insufficient to accurately represent both surface and bulk-like regions, potentially affecting the predicted behavior of surface magnetization. In metallic systems such as RuO2, the electronic wavefunctions of the top and bottom surfaces of the slab used in the calculation can penetrate deep into the bulk and mutually interact if the slab is not sufficiently thick, leading to an alteration of electronic properties of the true surface. Second, the origin of the surface magnetism was not comprehensively explored. More importantly, since their work focused on surface catalysis, it did not address how surface magnetization influences electronic and magnetic properties that are critical for spectroscopic and transport phenomena, which is highly relevant for spintronics applications [5, 25].
@@
In this paper, using first-principles calculations based on density functional theory (DFT), we show that surface magnetism spontaneously develops in RuO2(110). The spontaneous surface magnetization can be understood from the breaking of symmetry on the surface of RuO2(110), leading to a significant reconstruction of the electronic structure of RuO6 octahedra at the surface layer, i.e., in the band filling of the Ru $4d$ orbitals, resulting in sizable local moments. Remarkably, the presence of spontaneous surface magnetization leads to spin-polarized surface states and spin-dependent transport effects. These results can be important for understanding the experimental data on RuO2 obtained using spectroscopic and transport measurements such as spin- and angle-resolved photoemission spectroscopy (S-ARPES) [10, 26], spin-polarized scanning tunneling microscopy/spectroscopy (SP-STM/SP-STS), anomalous Hall measurements [3], and interfacial spin-dependent transport phenomena [27].
@@ II Computational Method
Bulk RuO2 was simulated using a tetragonal unit cell with a rutile structure ($P4_{2}/mnm$, SG 136). Due to the metallic property of bulk RuO2, a dense $k$-point mesh of 12×12×18 was used to sample the bulk BZ. The lattice constant was adopted from a recent experimental determination [36]. RuO2 films along [110] were built by reorienting the lattice vectors of the primitive cell, i.e., primitive lattice vectors along [001], [$01\overline{1}$ 0], and [110] become [100], [010], and [001] of the RuO2(110) surface unit cell, respectively, thus giving us a nonpolar stoichiometric slab. The symmetry of the bulk unit cell transitions to $Pmmm$ (SG 47) for stoichiometric slabs with an odd number of Ru layers. A $k$-point mesh of 12×6×1 was used for structural optimization of the surface unit cell, and a denser mesh of 18×9×1 was used for density of states (DOS) calculations. To eliminate spurious image interactions between slabs, a vacuum space of at least 20 Å was added to the slab normal direction. Internal coordinates of the slab were fully relaxed until the Hellman-Feynman force on each atom was smaller than 0.005 eV/Å  while keeping the lattice vectors fixed to the experimental bulk values. Due to the recent experimental observation of a non-magnetic ground state in bulk RuO2 [9, 10, 11, 12, 13, 14], we employed the PBE+$U$ approach with $U=0$ [37, 38, 39] to describe its electronic structure. Additionally, since spin-orbit coupling does not significantly affect the electronic and magnetic structures of RuO2, it was neglected in our calculations.
@@
To gain insights into the chemical bonding between atomic pairs contributing to the emergence of spin-polarized states at the surface of RuO2, we performed a crystal orbital Hamilton population (COHP) analysis [40, 41]. The COHP method decomposes the one-particle band energies into interactions between atomic orbitals of adjacent atoms. It effectively weighs the DOS by the corresponding Hamiltonian matrix elements, thus recovering the phase information of calculated wavefunctions that is otherwise missing in the band structure or DOS descriptions of material electronic structures. This enables the identification of bonding, non-bonding, and antibonding characteristics among the pairwise atoms. The stability of these interactions is quantified by the COHP values. A positive COHP value indicates bonding interactions, while a negative value represents an antibonding character. Traditionally, COHP analysis was done directly from atom-centered basis set calculations. However, since our calculations were done using the plane-wave basis set, a projection scheme from this basis set to a local orbital basis set was employed using the LOBSTER code [40, 41, 42]. In this projection scheme, the equivalence of the traditional COHP quantity is --pCOHP (negative of projected COHP). Our study applied COHP analysis to examine the bonding character and stability of the nearest-neighboring atomic pairs around Ru atoms on the surface and in bulk-like regions. For the projection, we employed the local basis function as defined in pbeVaspFit2015.
@@
Since spin-polarized scanning tunneling microscopy (SP-STM) can be used to recognize spin polarization effects on the surface [43], we simulated spin-resolved STM images employing the Tersoff-Hamann approximation [44]. In this approach, the tunneling current at the simulated probe tip (i.e., at a particular distance from the surface) in an STM experiment is proportional to the local density of states (LDOS) of the integrated electronic states ranging from the Fermi level to a predefined energy level given by a bias voltage $V$, roughly corresponding to an experimental voltage. The LDOS is given by:
$n(r,E)=\sum_{i}\left|\psi_{i}(r)\right|^{2}\delta\left(\varepsilon_{i}-E\right),$ (1)
and the tunneling current can be expressed as
@@
where $n(r,E)$ represents the LDOS at a given position $r$ and energy $E$. The LDOS can be evaluated from partial charge densities calculated from a pre-converged wavefunction. The partial charge density file from VASP was read by the HIVE-STM program [45], and STM images were generated using the constant-height method (at a height of $\sim 3$ above the highest atoms of the slab—the oxygen bridging atoms). For a given bias $V$, the simulated STM images reflect the contrast in the partial charge densities within the energy range $0<E-E_{\text{F}}<V$ (for $V>0$, positive bias) and $V<E-E_{\text{F}}<0$ (for $V<0$ for negative bias).
@@ III.1 Electronic and magnetic properties of bulk RuO2
Figure 1: Rutile-type crystal structure, electronic structure and chemical bonding analysis of bulk RuO2. (a) Rutile structure of RuO2 with two Ru sublattices in red and blue, oxygen in silver color, (b) Bulk Brillouin zone (BZ) shown in black and rotated so that [110] crystallographic direction normal to the (110) surface whose BZ shown in red, (c) spin-polarized electronic band structure of the bulk exhibiting no spin-polarization ($E_{\text{F}}$ set to 0), (d) projected density of states on Ru and O in the bulk region, and (e) the corresponding --pCOHP curve on the same energy scale.
The spin-resolved electronic band structure of the bulk is shown in Fig. 1(c), which does not exhibit spin polarization. In the assumed altermagnetic state, the electronic structure of interest lying along the $\Gamma$$\mathrm{M}$ high-symmetry path of the bulk Brillouin zone (BZ) shows significant spin splitting ($>1$ eV) as described by theoretical calculations of the band structure with relatively large on-site Hubbard $U$ values ($U\sim 1.5-2$ eV) [2, 7]. However, this spin splitting does not appear in the non-magnetic state. Along this momentum direction, three bands are present near the Fermi level, with an additional flat band lying around $-1.5$ eV below the Fermi level. When considering the electronic structure of the (110) slab, the $\Gamma$$\mathrm{M}$ path of the bulk BZ is of particular importance since bulk band dispersions along this path and parallel to it are projected onto the high-symmetry path $\overline{\Gamma}$¯¯$\overline{\mathrm{Y}}$ of the surface BZ, as illustrated in Fig. 1(b).
The DOS and chemical bonding analysis, shown in Fig. 1(d) and  1(e), highlights important characteristics of Ru–O bonds in RuO2. The DOS in Fig. 1(d) clearly shows strong hybridization between the Ru and O orbitals over a wide range of energy. Notably, between 8 and 6 eV below the Fermi level, pronounced peaks in the atomic-projected density of states (PDOS) and --pCOHP (positive—bonding character) with similar contributions from both Ru and O indicate a strong covalent bonding, best described by the $\sigma$-bonding composed of Ru $4d$ ($e_{g}$) and O $2p$ orbitals [46]. In addition, approximately from 6 eV to 3 eV below the Fermi level, small positive values of --pCOHP, with a greater contribution of oxygen, suggest a weaker covalent bonding character of the Ru–O bonds due to the $\pi$-bonding between Ru $4d$ ($t_{2g}$) and O $2p$. Around $-2$ eV, the contribution of Ru disappears, i.e., Ru PDOS is zero, leaving only O $2p$ states. This indicates that the electrons occupying these states belong to non-bonding orbitals and do not contribute to the Ru–O bonding strength, leading the --pCOHP values to approach zero.
@@
Despite the non-magnetic character of bulk RuO2, intriguing phenomena can arise at the (110) surface, leading to the emergence of surface magnetism. This disparity between the bulk and surface properties can be attributed to the altered electronic structure and bonding characteristics at the surface, which disrupt the symmetry and electronic interactions present in the bulk. The unique environment at the surface can facilitate localized magnetic moments that do not exist in the bulk material. In the next section, we will explore how these localized states, influenced by surface geometry and reduced coordination, contribute to the overall magnetic behavior of the (110) surface.
III.2 Magnetization on RuO2(110) surface
III.2.1 Emergence of surface magnetization